OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)?

OpenStudy (pfenn1):

I'm not sure what you are asking here. Do you need the equation for the line that passes through these two points?

OpenStudy (anonymous):

apply formula of m=y2-y1/x2-x1 and it will give you the value of slop that is called m. m = 3/4 . apply formula , y=mx+ b ,use values of x and y from one of the given values , i used ( 4 , 5) , so anwser comes out , b = 2 , now apply x= 0 to find out y intercept so y=mx+b ,y= 2x0 +2 , so y intercept is 2